I am not understanding this question my nephew has. He needs to Graph each function.

y= 2/x+6

Answer by LinnW(1048)

This is the same as y= (1/2)x+6
To do a graph try several values for x
When x = 0 , y = 6 so one point on the line is (0,6)
Try x = 2. y = (1/2)(2) + 6 = 1 + 6 = 7
So another point is (2,7)
With a strait edge draw a line through the two points.
Of course you can try additional values for x
and see what you get for y.